Shift report 20020107 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Jan 07 2002 - 08:14:44 EST

  • Next message: Apache: "Shift report 20020107 08:00-16:00"

    ------------------------------------------------------------
    Date: Mon Jan  7 08:14:44 2002, Shift: 00:00-08:00
    Supervisor: sjs, with: 
    Log book#12 pages 13 - 14
    
    Shift Summary:
    ------------------------------------------------------------
    1.) Continuous store throughout shift
    2.) Data runs:  (all 12 deg 1/8B; 40 deg 4A)
          run   effective   effective
                  trig 3       trig 6
          6318     244K          97K
          6319      91K          19K (see note below)
          6322      57K          12K
    
    With 392K effective trig 3 at this setting, called to have
    D5 polarity changed.  
    
    3.) D3 tripped , probably during run 6319 (was OK at
    beginning of run)
    
      Problems with HV on BBL and BBR supplies(discovered
    during BRAVO check at beginning of run 6319):
    Low Voltages (unable to increase): 
        BBR31-35, D1U1,D1D2,D1U3,D1P3,MRSL,MRSR
        BBL37-40
    Tripped (unable to reset):
        D1D1,D1U2   (FS trigger counters...)
        BBL41, BBL43, BBL44
    
      Checked timing of fera gate for Si detectors with beam:
    gate was about 500 ns early--adjusted...
    
    
     
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Jan 07 2002 - 08:15:25 EST